OpenStudy (anonymous):

A polynomial f(x) with real coefficients and leading coefficient 1 has zeros 2, -3 - 9i and degree 3. Express f(x) as a product of linear and quadratic polynomials with real coefficients that are irreducible over IR.

OpenStudy (anonymous):

f(x) = (x-2)(x+3+9i)(x+3-9i)

OpenStudy (anonymous):

then you can expand the last two

OpenStudy (anonymous):

How do you foil it to remove the i's?

OpenStudy (anonymous):

\[(a+z)(a+z conjugate ) \] = a^2 +a(z+z conj ) + z (z conj )

OpenStudy (anonymous):

cant believe they dont have the bars on top :|

OpenStudy (anonymous):

let a=x , z= 3+9i

OpenStudy (anonymous):

x^2 + x ( 3+9i +3-9i) + (3+9i)(3-9i)

OpenStudy (anonymous):

that should be it hopefully

OpenStudy (anonymous):

(x^2 +6x + (9 - (-81) )

OpenStudy (anonymous):

f(x) = (x-2)(x^2 +6x +90)

OpenStudy (anonymous):

fair sure thats it

OpenStudy (anonymous):

you can use quadratic formula to see if its correct.

OpenStudy (anonymous):

see if that gives the two complex roots

OpenStudy (anonymous):

and thats correct

OpenStudy (anonymous):

Wow thanks heaps!